OpenStudy (anonymous):

how do i multiply a 2 digit number with a 2 digit for example 36*47

OpenStudy (anonymous):

Lots of ways. The textbook method involves columns: 47 x 36 ==== 282 <--- 6 times 47 + 1410 ---- ^ 3 times 47 ^ | bring down 0 ======== 1692

OpenStudy (anonymous):

Personally I hate this way.

OpenStudy (anonymous):

I find it far easier to break up and distribute: \[47 \cdot 36 \]\[= (40 + 7)(30 + 6)\]\[=1200 + 210 + 240 + 42\]\[=1692\]